OpenStudy (anonymous):

Which of the following is not a geometric sequence? A. 1, 2, 4, 8 B. 3, 1, -1,-3 C. ½, ¼, 1/8, 1/16 D. 216, 72, 24, 8

OpenStudy (anonymous):

a

OpenStudy (zzr0ck3r):

in a geometric sequence \[\frac{2nd-term}{1st-term}=\frac{3rd-term}{2nd-term}......\]

OpenStudy (goformit100):

@codeblue11 don't give the answers directly

OpenStudy (anonymous):

@goformit100 ok sorry

OpenStudy (zzr0ck3r):

and your wrong @codeblue11

OpenStudy (zzr0ck3r):

(2/1) = (4/2) = (8/4) = 2

OpenStudy (anonymous):

So I should just follow that pattern with all the possible choices until one doesn't work out?

OpenStudy (zzr0ck3r):

yep we know 1 is geometric now look at 2 does (1/3) = (-1/1) = (-3/-1)?

OpenStudy (anonymous):

No, I don't think so.

OpenStudy (zzr0ck3r):

you don't think? does (1/3) = (-1/1)?

OpenStudy (anonymous):

1/3 is .33 and -1/1 is -1.

OpenStudy (anonymous):

So then it would be B.

OpenStudy (mathstudent55):

correct